OpenStudy (anonymous):

write a polynomial function of least degree with integral coefficients that have this given zero-> 3i

OpenStudy (anonymous):

if one zero is 3i then the conjugate of 3i that is -3i is also one of the zero.

OpenStudy (anonymous):

so the required polynomial is (x+3i)*(x-3i) =x^2 - 9i^2 =x^2 + 9 [ as i^2=-1]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

how bout if -5+i is the zero. what would be polynomial function for that if the conjugate is -5-i?